OpenStudy (anonymous):

MEDAL AND FAN! Which of the following statements best describes the graph of -5x+2y=1? A) It is a curve joining the points (-5,2),(2,3), and (4,1) B) It is a curve joining the points(-1,-3),(-1,-3), and (1,5) C) It is a straight line joining the points (1,3), (3,8), and (-3,-7) D) It is a straight line joining the points (4,-3), (-1,2), and (-4,5)

OpenStudy (anonymous):

@Michele_Laino do you know this one?

OpenStudy (anonymous):

@hartnn ?

OpenStudy (michele_laino):

I think is C. because if I substitute the coordinates of points (1,3), (3,8), and (-3,-7) in your equation, I got: \[-5*1+2*3=1, checked!\] \[-5*3+2*8=1, checked!\] and: \[-5*(-3)+2*(-7)=1, checked!\]

OpenStudy (anonymous):

:D thank you :3

OpenStudy (michele_laino):

thank you!!!!!